OpenStudy (anonymous):

Please help! I will fan+testimony+25 medals Explain in at least 5 sentences how to find the circumference of a circle, a semicircle, and a quarter circle. How is finding circumference of a quarter-circle different than finding the perimeter of a quarter circle?

OpenStudy (anonymous):

for a circle is pie*radius^2 for half cicle its ""/2 for quarter circle ""/4

OpenStudy (anonymous):

because it is a half of a circle guess if you have $2 what is the half/ you divide it by 2

OpenStudy (anonymous):

and think if you have $8 what woul e the quarter 8/4 right? thats how you do it

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

Thanks! your medals will be here in 15 minutes my dad wants me to help him with the car

OpenStudy (curiousquestions):

To get the circumference of a circle, you must know the radius or the diameter of the circle. If you know the radius of the circle, you must do the following equation: \[C=\Pi r ^{2}\] If you know the diameter, you just need to do this: \[C=\Pi d\] Now, to get the circumference of a simi-circle you need to divide the circumference of a whole circle by 2. If you need the circumference of a quarter-circle, you just divide the circumference of a whole circle by 4. The perimeter is different than the circumference because the perimeter is the distance around a polygon, whereas the circumference is the distance around a circle.

OpenStudy (curiousquestions):

That is 6 sentences. :)

OpenStudy (triciaal):

wait a minute you wrote the formula for the area

OpenStudy (triciaal):

C = 2*pi* r or C = pi* d since 2r = d The perimeter is the distance all around the polygon. the circumference is the perimeter of the special polygon the circle.